# SocialEyes: Scaling mobile eye-tracking to multi-person social settings

### v3

@@ A.1.1. Apparatus
The validation procedure was conducted with the existing Pupil Labs Neon eye-trackers in the same physical space as the Utility Test in 4. Centralview was captured with an ArduCam EK035 camera, featuring a global shutter and frame-wise hardware timestamping. An ArUco marker–25.5 inches wide and 21 inches tall–was mounted on a sturdy stand and displayed to the audience at 9 different positions on the stage. Neon devices were wirelessly connected to the local network and the current SocialEyes implementation was used for data collection and analysis.
@@ A.1.2. Procedure
19 participants were recruited during a public event at LIVELab, McMaster University. Participants seated themselves randomly among the audience seats and were equipped with the eye-trackers before the event started. The seating rows began at a distance of 134 inches from the stage front, with 50 inches of gap between each consecutive row. The validation procedure was conducted at the end of the hour-long event and was completed in roughly 3 minutes. During the procedure, an experimenter serially placed the ArUco marker (Figure 13B) at each of the positions– to –presented in Figure 13A. Participants were instructed to look at the ArUco target, ensuring that the target was visible in their egoviews.
@@ A.1.3. Data Analysis
100 consecutive frames from the centralview video were selected for each of the 9 presentation locations. ArUco markers were detected in time-synchronised egoviews of each participant and transformed to centralview using SocialEyes’ Homography module (Section 3.2.6). Projection errors were calculated as the Euclidean distances between the transformed and the detected marker centers in the centralview (640x480 resolution). Data from five devices were excluded due to more than 40% missing projections, either due to failure in marker detection or due to inadequate number of features detected for homography.
@@ A.3. Discussion
In summary, despite the added sensor noise in centralview data, the homogarphy planar assumption remains valid. The projection errors did not show significant variation across different scene depths. For future applications, it is important to ensure sufficient distance between the scene and observers for the assumption to hold, as projection errors may increase significantly at shorter distances.

### v4

@@ A.1.1. Apparatus
The validation procedure was conducted with the existing Pupil Labs Neon eye-trackers in the same physical space as the Utility Test (Section 4 of the main paper). Centralview was captured with an ArduCam EK035 camera, featuring a global shutter and frame-wise hardware timestamping. An ArUco marker–25.5 inches wide and 21 inches tall–was mounted on a sturdy stand and displayed to the audience at 9 different positions on the stage. Neon devices were wirelessly connected to the local network and the current SocialEyes implementation was used for data collection and analysis.
@@ A.1.2. Procedure
19 participants were recruited during a public event at LIVELab, McMaster University. Participants seated themselves randomly among the audience seats and were equipped with the eye-trackers before the event started. The seating rows began at a distance of 134 inches from the stage front, with 50 inches of gap between each consecutive row. The validation procedure was conducted at the end of the hour-long event and was completed in roughly 3 minutes. During the procedure, an experimenter serially placed the ArUco marker (Figure 13B) at each of the positions– to –presented in Figure 13A. Participants were instructed to look at the ArUco target while we counted down from five. Doing so ensured that the target was visible in their egoviews.
@@ A.1.3. Data Analysis
100 consecutive frames from the centralview video were selected for each of the 9 presentation locations. ArUco markers were detected in time-synchronised egoviews of each participant and transformed to centralview using SocialEyes’ Homography module (Section 3.2.6). Projection errors were calculated as the Euclidean distances between the transformed and the detected marker centers in the centralview (640x480 resolution). Data from five devices were excluded due to more than 40% missing projections, either due to failure in marker detection or due to an inadequate number of features detected for homography.
@@ A.3. Discussion
In summary, despite the added sensor noise in centralview data, the homography planar assumption remains valid. The projection errors did not show significant variation across different scene depths. For future applications, it is important to ensure sufficient distance between the scene and observers for the assumption to hold, as projection errors may increase significantly at shorter distances.